A total of 600 tickets were sold for a concert. Twice as many tickets were sold in advance than at the door. If the tickets sold in advance were $25 each and the tickets sold at the door were $32 each, how much money was collected for the concert?
Answer by Paul(988)

2x+x=600
3x=600
x=200
Now just plug in the values and find the money.
Advance = 25(2x)
Door = 32(x)